Session: Nov 1999 (session 4)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K.Borkowski at Sun Nov 28 11:41:04 MET 1999) Success
No known problems, except the observations started 1 minute late because of the FS time jump (corrected 5 min. later).

Shanghai (Huang Xinyong at Mon Nov 29 00:48:30 MET 1999) Success

Effelsberg (A. Kraus at Mon Nov 29 09:30:01 MET 1999) Minor failures
Several parity errors.

Medicina (G.Maccaferri at Mon Nov 29 10:48:38 MET 1999) Failure
Experiment suspended on doy 331-0030 UT for general power failure at antenna vertex. The problem was not brefly solvable. The operator has requested technical staff assistance. Last three tapes NOT recorded.

Noto (C.Trigilio at Mon Nov 29 11:11:00 MET 1999) Minor failures
Lost 6 scans from 4:17 to 5:01 (day 331) due to recorder problems at tape change. Raining most of the experiment.

Westerbork (Tony Foley at Mon Nov 29 15:49:27 MET 1999) Minor failures
Lost from 05:02 to 05:20 with array problem (someone set year 2000 software running for these scans). Mostly observed with 13 telescopes in the array.

Jodrell (Alastair Gunn at Wed Dec 1 03:40:32 MET 1999) Minor failures
Telescope parked due to high winds until about 16:20 UT. First two tapes are empty and therefore not sent to correlator. No other known problems.

HartRAO (Jonathan Quick at Thu Dec 2 11:40:32 MET 1999) Success
No problems

Onsala (John Conway at Mon Dec 13 09:56:14 MET 1999) Success
No known problems

Calibration (Jean\-Francois Desmurs at Wed Dec 15 13:18:15 MET 1999) Minor failures
The ANTAB formatted file for this experiment can be found on VLBEER in the nov99 directory under the name: es034b.antab. Read the comments in the file to get further information. (Rem : Jb bad Tsys values.)